# Towards a Science of AI Agent Reliability

### v1

@@ Abstract
AI agents are increasingly deployed to execute important tasks. While rising accuracy scores on standard benchmarks suggest rapid progress, many agents still continue to fail in practice. This discrepancy highlights a fundamental limitation of current evaluations: compressing agent behavior into a single success metric obscures critical operational flaws. Notably, it ignores whether agents behave consistently across runs, withstand perturbations, fail predictably, or have bounded error severity. Grounded in safety-critical engineering, we provide a holistic performance profile by proposing twelve concrete metrics that decompose agent reliability along four key dimensions: consistency, robustness, predictability, and safety. Evaluating 14 agentic models across two complementary benchmarks, we find that recent capability gains have only yielded small improvements in reliability. By exposing these persistent limitations, our metrics complement traditional evaluations while offering tools for reasoning about how agents perform, degrade, and fail.
@@ 1  Introduction
In July 2025, Replit’s AI coding assistant deleted an entire production database despite explicit instructions forbidding such changes [8, 55]. Months earlier, Washington Post columnist Geoffrey Fowler asked OpenAI’s Operator to find “cheap eggs” for delivery, only to find that the agent made an unauthorized \$31.43 purchase from Instacart, violating the company’s user confirmation safeguard before purchases [19, 42]. In 2024, the New York City government launched a chatbot for business assistance which consistently provided illegal advice—telling landlords they did not need to accept Section 8 housing vouchers—and gave different (incorrect) answers to ten journalists asking the same question [32]. In each case, agents that were judged to be reasonably capable by internal assessments displayed unreliable performance in real-world deployment, leading to costly failures. These evident shortcomings in reliability therefore raise a fundamental question:
@@
The dominant paradigm for agent evaluation offers little help in answering this question. Current practice centers on reporting mean task success rates [71, 25, 37, 39]: the percentage of benchmark tasks an agent completes correctly. This approach has clear advantages: it is simple to compute, easy to compare across systems, and provides a clean optimization target. But it obscures the behavioral properties that matter most for deployment. Accuracy cannot distinguish an agent that fails on a fixed, identifiable subset of tasks from one that fails unpredictably with the same rate. Yet the former permits systematic debugging while the latter does not. Accuracy also cannot distinguish benign failures (incomplete outputs, formatting errors) from catastrophic ones (deleted files, unauthorized actions), even though no practitioner would view them as interchangeable. Standard benchmarks do not report sensitivity to input perturbations, nor do they assess whether agents can recognize when they are likely to fail and abstain from returning bad predictions.
@@
AI agents face challenges that are similar in their scope and consequence as many of these domains: stochastic behavior, sensitivity to input variation, and the potential for high-cost failures when deployed in real-world settings. Yet, agent evaluations lack comparable mechanisms for understanding and improving their reliability. We address this gap by adapting the safety-critical perspective to agent evaluations, decomposing reliability into four dimensions (see Table 1): consistency (repeatable behavior across runs), robustness (stability under input and environmental perturbations), predictability (calibrated confidence and discrimination of correct/incorrect predictions), and safety (bounded severity when failures occur). Each dimension captures a property that matters for deployment but cannot be measured by accuracy alone. Across these dimensions, we propose a total of twelve concrete metrics that are independent of raw accuracy (see Section 3), enabling comparison of reliability across agents with different capability levels.
Applying these metrics to 14 agentic models across two benchmarks, we find that reliability gains lag noticeably behind capability progress (see Figure 1): despite steady accuracy improvements over 18 months of model releases, reliability only shows modest overall improvement. To analyze this widening gap, our paper provides two key contributions:
@@
    A formal taxonomy and metric suite: We translate qualitative safety-critical principles into computable metrics, enabling evaluation of agent reliability independently of task success.
@@
    A comprehensive reliability profile of modern agents: A detailed mapping of where state-of-the-art agentic models succeed and fail, isolating consistency and predictability as the dimensions requiring immediate research focus.
@@
This paper treats reliability as an empirically measurable property of agent behavior. We do not propose algorithms for improving reliability, though our metrics inform such efforts. Our perturbations model natural variation and incidental faults, not adversarial attacks or worst-case threat models. We also do not address broader questions of value alignment, though reliability as we define it is a precondition for safe deployment. Our reliability decomposition complements the AI safety taxonomy of Qi et al. [45]: while their safety definition spans alignment, honesty, and interpretability alongside operational concerns, we focus on the operational dimensions and contribute consistency (i.e., run-to-run repeatability grounded in safety-critical engineering) together with concrete, computable metrics for each dimension. We hope that this paper fosters work in measuring and improving reliability by practitioners adopting agents, developers building new agents, and researchers investigating the underlying causes.
@@ 2  A Cross-Domain Perspective of Reliability
Every safety-critical field ties reliability to consequence-aware risk assessment. Nuclear safety computes not just failure frequencies but expected consequences: health effects, economic impacts, environmental damage. Process industry standards tie development rigor to the failure severity. Aviation targets catastrophic failure probabilities below one in a billion per flight hour. The unifying principle: not all failures are equal. ML evaluation of safety has largely focused on compliance with harmful requests [2] or specific harmful behaviors like sycophancy [43]. These are important concerns for model deployment and governance, but distinct from consequence-aware assessment of how agents fail at legitimate tasks.
@@
These four dimensions emerge consistently across safety-critical engineering. ML research has begun addressing aspects of each, though often as isolated concerns rather than components of a unified account of reliability. Our contribution is to synthesize this connection: by grounding agent evaluation in cross-domain reliability practice, we provide a principled decomposition that organizes scattered ML efforts and identifies gaps. Crucially, all four dimensions are independent of raw capability. A highly capable system can be unreliable [70], and a less capable system can be highly reliable within its operating envelope. This separation is essential: improving capability does not automatically improve reliability, and evaluating one does not suffice for evaluating the other.
@@ 3.3 Predictability ($\mathcal{R}_{\text{Pred}}$)
Calibration ($P_{\text{cal}}$) measures whether stated confidence levels match empirical success rates: an agent claiming 80% confidence should succeed roughly 80% of the time. Poor calibration, such as consistent overconfidence, leads users to trust outputs they should instead verify, while under-confidence can result in unnecessary deferral. Discrimination ($P_{\text{AUROC}}$) assesses whether confidence scores successfully separate successes from failures. An agent can be systematically overconfident yet still assign higher confidence to tasks it will complete correctly. The Brier score ($P_{\text{brier}}$) provides a proper scoring rule [^2] A proper scoring rule incentivizes reporting true beliefs; deviation from the true probability worsens the expected score. that jointly penalizes mis-calibration and poor discrimination, offering a holistic view of predictive quality.
@@ 4  Experiments
We present a case study of our reliability metrics on two common benchmarks with complementary goals. Our experiments address four central questions:
  1. 1.
    Disentanglement: Do agents with similar task accuracy exhibit meaningfully different reliability profiles, or do reliability metrics simply recapitulate capability?
  2. 2.
    Dimension gaps: Which reliability dimensions show the largest deficits in current agents, and how wide is the gap between the strongest and weakest dimensions?
  3. 3.
    Generalization: Do agent reliability profiles generalize across benchmarks with different task structures, or are they specific to particular domains and evaluation settings?
  4. 4.
    Scaling: Does model scale improve reliability?
@@ 4.1 Setup
We evaluate our reliability metrics on two benchmarks that offer complementary reliability challenges (see Appendix D.1 for details):
@@ 4.2 Main Results
Figure 1 reveals a noticeable disconnect between capability progress and reliability gains across agentic models. Despite 18 months of model development, overall reliability only shows small improvements over time. Notably, reliability improvements are disproportionate across evaluation scenarios: highly structured environments ($\tau$-bench) show moderate gains, while open-ended tasks (GAIA) show barely any improvement, even among the latest models. These findings underscore that improving raw task performance is insufficient for building dependable AI agents; reliability requires targeted attention beyond capability scaling alone. Detailed trend plots for individual reliability metrics can be found in Figure 7.
@@ 4.3 Connection to Real-World Failures
We further revisit the real-world failures from Section 1 (Replit agent, OpenAI Operator, NYC chatbot) to examine whether our metrics would have provided adequate early warning signals. Our analysis is shown in Table 3 and concludes that each of the vulnerabilities could have been identified prior to deployment through systematic evaluation using our reliability metrics.
@@ 7  Conclusion
In addition to the suggested extensions in Section 6, we highlight several other directions for future work. Natural extensions of our work include modeling how reliability evolves over extended sessions where errors compound, extending these metrics to multi-agent systems where failures propagate across agents, and optimizing agents directly for reliability dimensions rather than capability alone. More broadly, we envision reliability evaluation becoming a standard component of model and scaffold releases, supported by an interdisciplinary community that brings together AI researchers and reliability engineers from safety-critical domains. On the deployment side, developing online signals that predict reliability failures before they manifest would enable proactive intervention. We elaborate on a broader set of future research questions for building more reliable AI agents in Appendix C.
@@ D.2 Implementation
We evaluate 14 language models from three major providers, spanning release dates from April 2024 to December 2025. All experiments are conducted using the Holistic Agent Leaderboard (HAL) [29] evaluation harness [^4] https://github.com/princeton-pli/hal-harness, which provides standardized agent execution, automatic cost tracking via Weave  [^5] https://docs.wandb.ai/weave, and reproducible evaluation protocols.

### v2

@@ Abstract
AI agents are increasingly deployed to execute important tasks. While rising accuracy scores on standard benchmarks suggest rapid progress, many agents still continue to fail in practice. This discrepancy highlights a major limitation of current evaluations: focusing on a single metric is not enough to understand agent behavior. Notably, it ignores whether agents behave consistently across runs, withstand perturbations, fail predictably, or have bounded error severity. Grounded in safety-critical engineering, we provide a holistic performance profile consisting of twelve metrics that decompose agent reliability along four key dimensions: consistency, robustness, predictability, and safety. Evaluating 14 models across two complementary benchmarks, we find that recent capability gains have only yielded small improvements in reliability. By exposing these persistent limitations, our metrics complement traditional evaluations while offering tools for reasoning about how agents perform, degrade, and fail.
@@ 1  Introduction
In July 2025, Replit’s AI coding assistant deleted an entire production database despite explicit instructions forbidding such changes [8, 58]. Months earlier, Washington Post columnist Geoffrey Fowler asked OpenAI’s Operator to find “cheap eggs” for delivery, only to find that the agent made an unauthorized \$31.43 purchase from Instacart, violating the company’s user confirmation safeguard before purchases [20, 46]. In 2024, the New York City government launched a chatbot for business assistance which consistently provided illegal advice and gave different (incorrect) answers to ten journalists asking the same question [34]. In each case, agents were judged to be reasonably capable by internal assessments, but displayed unreliable performance in real-world deployment, leading to costly failures. This raises a fundamental question:
@@
The dominant paradigm for agent evaluation offers little help in answering this question. Despite recent efforts of providing a more holistic evaluation of agents [31] and the language models that drive them [37], standard evaluation practice centers on reporting mean task success rates [76, 27, 40, 42]: the percentage of benchmark tasks an agent completes correctly. This approach has clear advantages: it is simple to compute, easy to compare across systems, and provides a clean optimization target. But it obscures the behavioral properties that matter most for deployment. Accuracy cannot distinguish an agent that fails on a fixed, identifiable subset of tasks from one that fails unpredictably with the same rate. Yet the former permits a partitioning of tasks between humans and agents, while the latter does not. Accuracy also cannot distinguish benign failures (incomplete outputs, formatting errors) from catastrophic ones (deleted files, unauthorized actions), even though no practitioner would view them as interchangeable. Standard benchmarks do not report sensitivity to input perturbations, nor do they assess whether agents can recognize when they are likely to fail and abstain from returning bad predictions.
@@
We adapt the safety-critical perspective to agent evaluations, decomposing reliability into four dimensions (see Table 1): consistency (repeatable behavior across runs), robustness (stability under input and environmental perturbations), predictability (calibrated confidence and discrimination of correct/incorrect predictions), and safety (bounded severity when failures occur). Each dimension captures a property that matters for deployment but cannot be measured by accuracy alone. Across these dimensions, we propose a total of twelve concrete metrics that are independent of raw accuracy (see Section 3), enabling comparison of reliability across agents with different capability levels.
Applying these metrics to 14 models across two benchmarks, we find that reliability gains lag noticeably behind capability progress (see Figure 1): despite steady accuracy improvements over 18 months of model releases, reliability only shows modest overall improvement. To analyze this widening gap, our paper provides two key contributions:
@@
    A formal taxonomy and suite of metrics: We translate qualitative safety-critical principles into computable metrics, enabling evaluation of agent reliability independently of task success.
@@
    A comprehensive reliability profile of modern agents: A detailed mapping of where state-of-the-art models succeed and fail, isolating consistency and predictability as the dimensions requiring immediate research focus.
@@
We treat reliability as an empirically measurable property of agent behavior. We do not propose algorithms for improving reliability, though our metrics inform such efforts.
We view reliability as one of many properties beyond capability that are needed for AI agents to be useful and trustworthy. For example, cost, latency, transparency / explainability, collaboration ability, and value alignment could all be important, but we do not consider them. Most significantly, our perturbations model natural variation and incidental faults, not adversarial attacks or worst-case threat models.
For broader conceptions of safety, see  [15, 23, 67, 68, 45, 51]. We use the term “safety” narrowly and limit ourselves to operational concerns.
@@ 2  A Cross-Domain Perspective of Reliability
Every safety-critical field ties reliability to consequence-aware risk assessment. Nuclear safety computes not just failure frequencies but expected consequences: health effects, economic impacts, environmental damage. Process industry standards tie development rigor to the failure severity. Aviation targets catastrophic failure probabilities below one in a billion per flight hour. The unifying principle is that not all failures are equal. ML evaluation of safety has largely focused on compliance with harmful requests [2] or specific harmful behaviors like sycophancy [47]. These are important concerns for model deployment and governance, but distinct from consequence-aware assessment of how agents fail at legitimate tasks.
@@
These four dimensions emerge consistently across safety-critical engineering. ML research has begun addressing aspects of each, though often as isolated concerns rather than components of a unified account of reliability. In doing so, we take inspiration from Liang et al. [37], which provides a comprehensive multi-metric taxonomy on evaluation of language models. Our contribution is to synthesize this connection for AI agents: by grounding agent evaluation in cross-domain reliability practice, we provide a principled decomposition that organizes scattered ML efforts and identifies gaps. Crucially, all four dimensions are independent of raw capability. A highly capable system can be unreliable [75], and a less capable system can be highly reliable within its operating envelope. This separation is essential: improving capability does not automatically improve reliability, and evaluating one does not suffice for evaluating the other.
@@ 3.3 Predictability ($\mathcal{R}_{\text{Pred}}$)
Calibration ($P_{\text{cal}}$) measures whether stated confidence levels match empirical success rates: an agent claiming 80% confidence should succeed roughly 80% of the time. Poor calibration, such as consistent overconfidence, leads users to trust outputs they should instead verify, while under-confidence can result in unnecessary deferral. Discrimination ($P_{\text{AUROC}}$) assesses whether confidence scores successfully separate successes from failures. An agent that assigns 90% confidence to all outputs might be poorly calibrated, but if correct answers consistently receive higher scores than incorrect ones, the agent still discriminates well. Good discrimination enables users to set a confidence threshold above which outputs are accepted and below which they are deferred. The Brier score ($P_{\text{brier}}$) provides a proper scoring rule [^2] A proper scoring rule incentivizes reporting true beliefs; deviation from the true probability worsens the expected score. that jointly measures calibration and discrimination, offering a holistic view of predictive quality.
@@ 4  Experiments
We evaluate reliability on two established benchmarks that pose complementary challenges. Our goal is to understand the current state of agent reliability, both in aggregate and across individual sub-metrics, and how it has evolved over recent model generations.
@@ 4.1 Setup
We select the following two benchmarks for their complementary reliability challenges (see Appendix D.1 for details):
@@ 4.2 Main Results
Figure 1 reveals a noticeable disconnect between capability progress and reliability gains across models. Despite 18 months of model development, overall reliability only shows small improvements over time. Notably, reliability improvements are disproportionate across evaluation scenarios: $\tau$-bench shows moderate gains, while GAIA shows barely any improvement, even among latest models. One possible explanation is that the more structured nature of $\tau$-bench makes consistent behavior easier to achieve, whereas GAIA’s open-ended tasks present a wider range of failure modes; though other factors (e.g., benchmark difficulty, evaluation protocol) may also contribute. Regardless of the cause, these findings suggest that improving raw task performance may not be sufficient for building dependable AI agents, and that reliability may require targeted attention beyond capability scaling alone. Detailed trend plots for individual reliability metrics can be found in Figure 7.
@@
  • Safety: We report aggregate safety ($\mathcal{R}_{\text{Saf}}$) alongside its two sub-metrics: compliance ($S_{\text{comp}}$) and harm severity ($S_{\text{harm}}$). Recent frontier models exhibit markedly lower violation rates (see Figure 5), with the most capable models from each provider achieving the highest compliance scores. Harm severity scores are generally high across the board, indicating that when violations do occur, most are low-to-moderate in severity. However, even infrequent high-severity violations—such as unauthorized data exposure or incorrect financial transactions—carry outsized costs. We provide a detailed breakdown of safety on $\tau$-bench below.
@@ 4.3 Connection to Real-World Failures
We further revisit the real-world failures from Section 1 (Replit agent, OpenAI Operator, NYC chatbot) to examine whether our metrics would have provided early warning signals. Our analysis in Table 3 concludes that each of the vulnerabilities could have been identified prior to deployment through systematic evaluation using our reliability metrics.
@@ 6  Limitations
  • Choice of temperature. Where applicable, we set the temperature to zero in all our experiments. This limits one major source of stochasticity in model outputs. When attempting to maximize the accuracy of an agent, a nonzero temperature might be required, and our experiments might overestimate the reliability that is achievable in such circumstances.
@@ 7  Conclusion
In addition to the suggested extensions in Section 6, we highlight several other directions for future work. Natural extensions of our work include modeling how reliability evolves over extended sessions where errors compound, extending these metrics to multi-agent systems where failures propagate across agents, and optimizing agents directly for reliability dimensions rather than capability alone. More broadly, we envision reliability evaluation becoming a standard component of model and scaffold releases, supported by an interdisciplinary community that brings together AI researchers and reliability engineers from safety-critical domains. On the deployment side, developing online signals that predict reliability failures before they manifest would enable proactive intervention. Appendix C elaborates on a broader set of future research questions.
@@ D.2 Implementation
We evaluate 14 language models from three major providers, spanning release dates from April 2024 to December 2025. All experiments are conducted using the Holistic Agent Leaderboard (HAL) [31] evaluation harness [^4] https://github.com/princeton-pli/hal-harness, which provides standardized agent execution, automatic cost tracking via Weave  [^5] https://docs.wandb.ai/weave, and reproducible evaluation protocols. We note that for OpenAI models we use the Chat Completions API instead of the Responses API [^6] https://developers.openai.com/api/docs/guides/migrate-to-responses/.